THE LINPACK BENCHMARK ON THE AP1000

RICHARD P. BRENT

Abstract

We describe an implementation of the Linpack Benchmark on the Fujitsu AP 1000. Design considerations include communication primitives, data distribution, use of blocking to reduce memory references, and effective use of the cache. Some details of our blocking strategy appear to be new. The results show that the Linpack Benchmark can be implemented efficiently on the AP 1000, and it is possible to attain about 80 percent of the theoretical peak performance for large problems.

Comments

Only the Abstract is given here. The full paper appeared as [2]. For related work see [3].

References

1. R. P. Brent (editor), Proceedings of the Second Fujitsu-ANU CAP Workshop, Department of Computer Science, Australian National University, November 1991, 254 pp. rpb129.

2. R. P. Brent, “The Linpack benchmark on the AP 1000”, Proc. Frontiers ’92 (McLean, Virginia, October 1992), IEEE Press, 1992, 128–135. ISBN 0-8186-2772-7. A preliminary report appeared in [1, G1–G13]. rpb130.

3. R. P. Brent and P. E. Strazdins, “Implementation of the BLAS level 3 and Linpack benchmark on the AP 1000”, Fujitsu Scientific and Technical Journal 29, 1 (March 1993), 61–70. rpb136.

4. P. E. Strazdins and R. P. Brent, “The implementation of BLAS level 3 on the AP 1000: Preliminary report”, in [1, H1–H17]. rpb131.

Computer Sciences Laboratory, Australian National University, Canberra E-mail address: rpb@cs.anu.edu.au

1991 Mathematics Subject Classification. Primary 65Y05; Secondary 65F05, 65Y10.

Key words and phrases. BLAS, basic linear algebra subroutines, Linpack benchmark, parallel computation, distributed memory, local memory, MIMD, Fujitsu AP1000, Gaussian elimination, block methods, partial pivoting.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1992, IEEE.

Comments \begin{matrix}c \\ ◯ \end{matrix} 1993, R. P. Brent. rpb130a typeset using AMS-L^{A}TEX